OpenStudy (anonymous):

what is an equation of the line that is perpendicular to 2y+2= -7/4(x-7) and contains (-2,-3)

OpenStudy (anonymous):

first solve 2y+2=-7/4(x-7) \[ y = -\frac{7x}{8}+\frac{41}{8}\] perpendicular line's slope = -1/m => 8/7 y-(-2)=8/7(x-(-3)) => \[y=\frac{8}{7}x+\frac{10}{7}\]

OpenStudy (anonymous):

Thanks when you get done i know the answer is y=8/7x-5/7.